# PneuGelSight: Soft Robotic Vision-Based Proprioception and Tactile Sensing

### v1

@@ 1 Introduction
Our previous works Yoo et al. (2023); Wang et al. (2020a) first introduced the method of using internal vision and deep learning techniques to reconstruct the high-dimensional 3D shapes of the soft robots. The proprioception sensing capability of PneuGelSight extends this prior work by improving both the visual pattern design and algorithms. These enhancements enable the sensor to support both proprioception and tactile sensing using a single onboard camera, while also delivering significantly improved performance in shape reconstruction. Experiments demonstrate that our system accurately senses both large-scale deformations due to actuation or environmental interaction and fine details of object surfaces in contact. This capability enables soft grippers to better understand both the global and local shape of objects, enhancing object recognition and property sensing. Our system significantly improves the ability of soft robots to perform complex grasping and manipulation tasks. The proposed methodology of using deep learning for high-dimensional modeling and simulation for sensor design can be extended to the sensing system design for other soft robots.
@@ 2.1 Sensing for Soft Robot
Alternatively, researchers have looked into novel solutions for soft robotic sensing in recent years, including acoustic (Wall et al. (2023)) and vision-based sensing(Yoo et al. (2023); Nakao et al. (1990)). Acoustic sensing has been shown to work well with soft structures, achieving sensing with a simple embedded microphone. But like string sensors, the sound received at the end-effector is aggregated through the whole body, limiting the sensing resolution. Vision-based sensing, originally proposed for precise contact texture reconstruction, remains underexplored in soft robotics. Our previous work, Yoo et al. (2023) , demonstrated high-resolution perception in a soft robot by leveraging a hollow structure embedded with densely distributed internal visual markers. An internal camera captured the motion of these markers during deformation, and a simple neural network was used to reconstruct the robot’s real-time shape as a dense point cloud. That study also introduced a sim-to-real pipeline in zero seconds to streamline data collection for model training.
@@ 2.2 Vision-based Tactile Sensing
In contrast, our objective is to fully sensorize a soft robot with high-resolution proprioception and tactile sensing, which fully leveraging the potential of vision-based sensing principles exemplified by GelSight. While the core principle remains similar, applying it to a highly deformable robot introduces fundamental challenges. Traditional algorithms for shape and contact reconstruction fail, as they cannot accommodate non-planar deformations where the entire sensor geometry evolves dynamically. Moreover, optical signals are compromised under strain due to variations in illumination, reflection, and surface normal estimation.
@@ 3.2 Design and Fabrication
We 3D print the actuated body of PneuGelSight with a compliant and flexible silicone material (Silicone 40A). The lighting system of the robot is configured into a distinct pattern: 6 green, 7 blue, 7 red, and 4 additional blue fibers, arranged in clockwise order. We design some semi-cylindrical protrusion on the side of the robot backbone to constrain the orientation of the fibers. Each of the fibers is glued and inserted into the protrusion manually. After the lighting system is established, we seal the embedded camera into the camera hole and apply more silicon glue for air tightness. The camera’s field of view is approximately 160 degrees (Arducam, Wide Angle Camera).
@@ 4 Proprioception: Sensing Robot Deformation
The machine learning model uses the finger’s geometric features as input. Specifically, it extracts the front surface contour from a binary image and outputs a feature describing the finger’s detailed deformation based on the original point-cloud shape. To improve the efficiency of data collection, we develop a physics-based simulation framework to collect the camera’s data under different robot deformation cases, and then perform a zero-shop transfer to the real robot. Our new method includes a PointNet(Qi et al. (2017)) auto-encoder to address the new challenge of the geometrical feature as input since the complicated nature of the sensor design contained our choice of customizing the geometrical markers to extract from the finger. Unlike the previous method, which directly predicts the finger’s shape, our new model learns a deformation feature that is applied to the original shape and efficiently captures high-dimensional deformation from simple input features by incorporating shape priors.
@@ 5.1 Region Proposal and Background Estimation Module
Figure 7: Effect of point could density on proprioception accuracy.
@@ 6.1 High-Resolution Proprioception
A demonstration of the proprioception result is shown in Figure 6. We quantify the results by measuring the standard chamfer distance in mm between the ground truth captured by RGBD camera and the predicted points, where lower values indicate better accuracy. The metric is calculated through the whole actuation process, and for simplicity, we only visualize the most significant scene in Figure 6.
We select five general test cases, two of which are neural pose and natural bending to represent typical scenarios during gripper operation, and others are bending forward, backward and lateral bending with external load. We present both side view (aligned with external image) and top views to help better interpret the point clouds. As visualized in Figure 7, our model achieves Chamfer Distances of 2.21, 6.76, 8.76, 2.12, and 7.09 mm for each test case, using a point cloud density of N=4096. The mean Chamfer Distance across all collected cases is 5.35 mm. Notably, even in the edge case of lateral bending—which is mechanically constrained and potentially harmful in real-world scenarios—the model demonstrates reasonable accuracy in reconstructing the deformation.
@@ 7 Evaluation of PneuGelSight in Real World Applications
We mount our soft gripper on a UR5e robot arm, and the avocado is lying flat on the table surface without any prior knowledge of its shape. At each time step, the last joint of the robot arm rotates at a pre-defined angle, and the gripper grips the avocado while capturing an image from inside the gripper. Each image provides an estimated bending shape of the gripper and a detailed contact texture reconstruction. Combining these results, we achieve a comprehensive shape and texture reconstruction of the avocado, as shown in the sensing result column of Figure 10. The reconstructed meshes are then stitched together for the final reconstruction. The predicted surface shows strong alignment with each other, and although the sensing area is currently focused on the middle part of the avocado due to the gripper’s partial use of its capacity, the overall sensing area offers significant potential.

### v2

@@ 1 Introduction
Our previous works (Yoo et al. (2023); Wang et al. (2020a)) first introduced the method of using internal vision and deep learning techniques to reconstruct the high-dimensional 3D shapes of the soft robots. The proprioception sensing capability of PneuGelSight extends this prior work by improving both the visual pattern design and algorithms. These enhancements enable the sensor to support both proprioception and tactile sensing using a single onboard camera, while also delivering significantly improved performance in shape reconstruction. Experiments demonstrate that our system accurately senses both large-scale deformations due to actuation or environmental interaction and fine details of object surfaces in contact. This capability enables soft grippers to better understand both the global and local shape of objects, enhancing object recognition and property sensing. Our system significantly improves the ability of soft robots to perform complex grasping and manipulation tasks. The proposed methodology of using deep learning for high-dimensional modeling and simulation for sensor design can be extended to the sensing system design for other soft robots.
@@ 2.1 Sensing for Soft Robot
Alternatively, researchers have looked into novel solutions for soft robotic sensing in recent years, including acoustic (Wall et al. (2023)) and vision-based sensing(Yoo et al. (2023); Nakao et al. (1990)). Acoustic sensing has been shown to work well with soft structures, achieving sensing with a simple embedded microphone. But like string sensors, the sound received at the end-effector is aggregated through the whole body, limiting the sensing resolution. Vision-based sensing, originally proposed for precise contact texture reconstruction, remains underexplored in soft robotics. Our previous work, Yoo et al. (2023) , demonstrated high-resolution perception in a soft robot by leveraging a hollow structure embedded with densely distributed internal visual markers. An internal camera captured the motion of these markers during deformation, and a simple neural network was used to reconstruct the robot’s real-time shape as a dense point cloud. That study also introduced a sim-to-real pipeline in zero-shot fashion to streamline data collection for model training.
@@ 2.2 Vision-based Tactile Sensing
In contrast, our objective is to fully sensorize a soft robot with high-resolution proprioception and tactile sensing, which fully leverages the potential of vision-based sensing principles exemplified by GelSight. While the core principle remains similar, applying it to a highly deformable robot introduces fundamental challenges. Traditional algorithms for shape and contact reconstruction fail, as they cannot accommodate non-planar deformations where the entire sensor geometry evolves dynamically. Moreover, optical signals are compromised under strain due to variations in illumination, reflection, and surface normal estimation.
@@ 3.2 Design and Fabrication
We 3D print the actuated body of PneuGelSight with a compliant and flexible silicone material (Silicone 40A). The lighting system of the robot is configured into a distinct pattern: 6 green, 7 blue, 7 red, and 4 additional blue fibers, arranged in clockwise order. We design some semi-cylindrical protrusion on the side of the robot backbone to constrain the orientation of the fibers. Each fiber is glued and inserted into the protrusion manually. After the lighting system is established, we seal the embedded camera into the camera hole and apply more silicon glue for air tightness. The camera’s field of view is approximately 160 degrees (Arducam, Wide Angle Camera).
@@ 4 Proprioception: Sensing Robot Deformation
The machine learning model uses the finger’s geometric features as input. Specifically, it extracts the front surface contour from a binary image and outputs a feature describing the finger’s detailed deformation based on the original point-cloud shape. To improve the efficiency of data collection, we develop a physics-based simulation framework to collect the camera’s data under different robot deformation cases, and then perform a zero-shot transfer to the real robot. Our new method includes a PointNet(Qi et al. (2017)) autoencoder to address the new challenge of the geometrical feature as input since the complicated nature of the sensor design contained our choice of customizing the geometrical markers to extract from the finger. Unlike the previous method, which directly predicts the finger’s shape, our new model learns a deformation feature that is applied to the original shape and efficiently captures high-dimensional deformation from simple input features by incorporating shape priors.
@@ 5.1 Region Proposal and Background Estimation Module
Figure 7: Effect of point cloud density on proprioception accuracy.
@@ 6.1 High-Resolution Proprioception
A demonstration of the proprioception result is shown in Figure 6. We quantify the results by measuring the standard chamfer distance in mm between the ground truth captured by RGB-D camera and the predicted points, where lower values indicate better accuracy. The metric is calculated through the whole actuation process, and for simplicity, we only visualize the most significant scene in Figure 6.
We select five general test cases, two of which are neutral pose and natural bending to represent typical scenarios during gripper operation, and others are bending forward, backward and lateral bending with external load. We present both side view (aligned with external image) and top views to help better interpret the point clouds. As visualized in Figure 7, our model achieves Chamfer Distances of 2.21, 6.76, 8.76, 2.12, and 7.09 mm for each test case, using a point cloud density of N=4096. The mean Chamfer Distance across all collected cases is 5.35 mm. Notably, even in the edge case of lateral bending—which is mechanically constrained and potentially harmful in real-world scenarios—the model demonstrates reasonable accuracy in reconstructing the deformation.
@@ 7 Evaluation of PneuGelSight in Real World Applications
We mount our soft gripper on a UR5e robot arm, and the avocado is lying flat on the table surface without any prior knowledge of its shape. At each time step, the last joint of the robot arm rotates at a predefined angle, and the gripper grips the avocado while capturing an image from inside the gripper. Each image provides an estimated bending shape of the gripper and a detailed contact texture reconstruction. Combining these results, we achieve a comprehensive shape and texture reconstruction of the avocado, as shown in the sensing result column of Figure 10. The reconstructed meshes are then stitched together for the final reconstruction. The predicted surface shows strong alignment with each other, and although the sensing area is currently focused on the middle part of the avocado due to the gripper’s partial use of its capacity, the overall sensing area offers significant potential.